# LLMSQL: Upgrading WikiSQL for the LLM Era of Text-to-SQL

### v1

@@ Abstract
Converting natural language questions into SQL queries (Text-to-SQL) enables non-expert users to interact with relational databases and has long been a central task for natural language interfaces to data. While the WikiSQL dataset played a key role in early NL2SQL research, its usage has declined due to structural and annotation issues, including case sensitivity inconsistencies, data type mismatches, syntax errors, and unanswered questions. We present LLMSQL, a systematic revision and transformation of WikiSQL designed for the LLM era. We classify these errors and implement automated methods for cleaning and re-annotation. To assess the impact of these improvements, we evaluated multiple large language models (LLMs), including Gemma 3, LLaMA 3.2, Mistral 7B, gpt-oss 20B, Phi-3.5 Mini, Qwen 2.5, OpenAI o4-mini, DeepSeek R1 and others. Rather than serving as an update, LLMSQL is introduced as an LLM-ready benchmark: unlike the original WikiSQL, tailored for pointer-network models selecting tokens from input, LLMSQL provides clean natural language questions and full SQL queries as plain text, enabling straightforward generation and evaluation for modern natural language-to-SQL models.
@@ I Introduction
Beyond WikiSQL, subsequent benchmarks broadened the scope of Text-to-SQL evaluation. Spider [1] emphasizes cross-domain generalization with multi-table queries, while BIRD [2] focuses on robustness to annotation noise in real-world settings. These efforts illustrate complementary directions, yet systematic cleaning of WikiSQL itself has remained underexplored despite its historical importance and widespread use.
@@ II Related Work
Converting natural language questions into SQL (NL2SQL) is a central problem for natural database interfaces, with progress driven by large annotated corpora and neural architectures. A key dataset is WikiSQL, released with Seq2SQL (Zhong et al. [3]), which introduced execution-based reinforcement learning to optimize queries by both text and results. Seq2SQL’s pointer-network design—selecting tokens from the input—limits its applicability to modern LLMs.
@@
Subsequent research focused on better capturing SQL structure and schema. SQLNet [4] used a sketch-based, slot-filling decoder for SELECT, aggregation, and WHERE components. TypeSQL [5] added type-aware features linking question tokens to table columns, improving handling of rare values. Coarse-to-Fine methods first predict a skeleton, then detailed WHERE conditions. Execution-aware approaches like SeaD [6] and execution-guided decoding [7] prune invalid queries, improving execution accuracy, which remains the main metric for WikiSQL.
@@
Execution-aware decoding techniques have also proven effective. Execution-guided decoding runs partial SQL against the target table during generation and prunes candidate continuations that lead to execution errors or empty/invalid results; this approach has been shown to reduce syntactic and semantic errors in generated queries and improve the share of successfully executable outputs [7].
@@
The integration of pretrained transformers with table-aware contextualization produced further gains beyond earlier seq2seq models. For instance, SQLova [8] adapts BERT-style encoders to incorporate table headers and cell values, yielding stronger logical-form and execution accuracy. Complementary approaches enhance schema representations and alignment between natural language and table content. These advances illustrate that pretrained models can surpass execution-guided seq2seq baselines. However, such results hold only for single-table WikiSQL and should be interpreted with caution when generalizing to multi-table or cross-domain settings
@@
Recognizing these limitations, new benchmarks were introduced. Spider [1] established a cross-domain standard with multi-table queries, JOINs, and nested structures, requiring compositional reasoning. UNITE [9] aggregated multiple datasets for broader SQL pattern coverage, while BIRD [2] emphasized robustness under noisy annotations. These resources expand scope, but none directly address the persistent quality issues in WikiSQL itself.
@@
A critical body of work addresses annotation quality, evaluation methodology, and robustness. Finegan-Dollak et al. [10] analyzed evaluation practices and demonstrated that certain dataset splits and anonymization choices can inflate reported performance; they proposed more conservative splitting strategies and evaluation protocols to better assess compositional generalization. Recent research, such as studies on BIRD-Bench, has highlighted that crowd-sourced datasets often contain issues such as noisy annotations, mismatched data types, and questions that cannot be answered from the table. These problems can bias standard evaluation metrics, including both execution accuracy and logical-form accuracy, giving a misleading impression of model performance [2]. To address this, researchers recommend cleaning datasets and performing explicit verification of model outputs. Additionally, broader benchmark analyses emphasize the need for standardized evaluation procedures and the inclusion of realistic, user-generated queries, which help ensure that models are tested on scenarios closer to real-world usage and that reported metrics more accurately reflect practical performance [11, 9].
@@
Taken together, the literature highlights three concurrent directions: (1) architectural and schema-aware improvements to increase generation correctness (e.g., type-aware models, table-aware encoding models, execution-guided decoding) [8, 5, 7]; (2) the design of more representative and cleaner benchmarks and evaluation practices that mitigate annotation noise (e.g., Spider, UNITE, and the recommendations of Finegan-Dollak) [1, 10, 9]; and (3) the systematic evaluation of large pretrained and/or instruction-tuned LMs for Text-to-SQL, along with the development of practical engineering techniques (prompting, constrained decoding, schema linking, execution validation) required to obtain executable and reliable queries in real-world settings [12, 13].
@@ III-A Introduction to NL2SQL and the WikiSQL Dataset
WikiSQL contains 80,654 (question, SQL) pairs derived from 24,241 Wikipedia tables [3]. It has become a standard benchmark due to its scale, simple query structures, and public availability.
@@ IV-1 Temperature
Recent studies have shown that using greedy decoding to evaluate reasoning models can lead to higher repetition rates and significant variability across different checkpoints [14]. Therefore, for models with strong reasoning capabilities, we did not use greedy decoding. Instead, we employed the default sampling parameters recommended for each model.
@@ IV-4 SQL Extraction from Outputs
To address this, we adopted a regex-based extraction strategy similar to what is used in popular text-to-SQL benchmarks such as Spider [1]. Specifically, we designed a regex pattern to extract all possible SQL statements from the model output:
@@ IV-A Main Scenario (Benchmark)
We evaluated a broad set of models across different scenarios, including Llama 3.2 1B Instruct, Qwen 2.5 1.5B Instruct, Phi 3.5 Mini Instruct, Gemma 3 4B IT, Gemma 3 27B IT, Mistral 7B Instruct v0.2, gpt-oss-20B, DeepSeek V3 0324, DeepSeek R1 0528, and OpenAI o4-mini [15, 16, 17, 18, 19, 20, 21, 14, 22].
@@ V Conclusions
We also argue that, despite its simplicity, LLMSQL remains a relevant and informative benchmark. Real-world SQL workloads are often dominated by relatively simple patterns: in an analysis of 8.1 million production queries at Uber [23], over 62% used JOIN, and fewer than 1% involved operators such as UNION, INTERSECT, or EXCEPT. This reinforces the idea that mastering simple clauses is critical for practical deployment and benchmarking progress.

### v2

@@ Abstract
Converting natural language questions into SQL queries (Text-to-SQL) enables non-expert users to interact with relational databases and has long been a central task for natural language interfaces to data. While the WikiSQL dataset played a key role in early NL2SQL research, its usage has declined due to structural and annotation issues, including case-sensitivity inconsistencies, data types mismatches, syntax errors, and unanswered questions. We present LLMSQL, a systematic revision and transformation of WikiSQL designed for the LLM era. We classify these errors and implement automated methods for cleaning and re-annotation. To assess the impact of these improvements, we evaluated multiple large language models (LLMs), including Gemma 3, LLaMA 3.2, Mistral 7B, gpt-oss 20B, Phi-3.5 Mini, Qwen 2.5, OpenAI o4-mini, DeepSeek-R1, and others. Notably, DeepSeek-R1 achieves 88.40% accuracy in a zero-shot setting, and models under 10B parameters surpass 90% accuracy after finetuning. Rather than serving as an update, LLMSQL is introduced as an LLM-ready benchmark: unlike the original WikiSQL—tailored for pointer-network models selecting tokens from input—LLMSQL provides clean natural-language questions and full SQL queries as plain text, enabling straightforward generation and evaluation for modern natural-language-to-SQL models.
@@ I Introduction
Beyond WikiSQL, subsequent benchmarks broadened the scope of Text-to-SQL evaluation. Spider [Spider2018] emphasizes cross-domain generalization with multi-table queries, while BIRD [Wretblad2024BIRD] focuses on robustness to annotation noise in real-world settings. These efforts illustrate complementary directions, yet systematic cleaning of WikiSQL itself has remained underexplored despite its historical importance and widespread use.
@@
    We release the official evaluation package for the LLMSQL benchmark [^1] https://github.com/LLMSQL/llmsql-benchmark, which provides standardized scoring, execution-based evaluation, and reference utilities for researchers and practitioners.
  2. 2.
@@ II Related Work
Converting natural language questions into SQL (NL2SQL) is a central problem for natural database interfaces, with progress driven by large annotated corpora and neural architectures. A key dataset is WikiSQL, released with Seq2SQL (Zhong et al. [Zhong2017Seq2SQL]), which introduced execution-based reinforcement learning to optimize queries by both text and results. Seq2SQL’s pointer-network design—selecting tokens from the input—limits its applicability to modern LLMs.
@@
Subsequent research focused on better capturing SQL structure and schema. SQLNet [Xu2017SQLNet] used a sketch-based, slot-filling decoder for SELECT, aggregation, and WHERE components. TypeSQL [Yu2018TypeSQL] added type-aware features linking question tokens to table columns, improving handling of rare values. Coarse-to-Fine methods first predict a skeleton, then detailed WHERE conditions. Execution-aware approaches like SeaD [Xu2021SeaD] and execution-guided decoding [Wang2018EG] prune invalid queries, improving execution accuracy, which remains the main metric for WikiSQL.
@@
Execution-aware decoding techniques have also proven effective. Execution-guided decoding runs partial SQL against the target table during generation and prunes candidate continuations that lead to execution errors or empty/invalid results; this approach has been shown to reduce syntactic and semantic errors in generated queries and improve the share of successfully executable outputs [Wang2018EG].
@@
The integration of pretrained transformers with table-aware contextualization produced further gains beyond earlier seq2seq models. For instance, SQLova [Hwang2019SQLova] adapts BERT-style encoders to incorporate table headers and cell values, yielding stronger logical-form and execution accuracy. Complementary approaches enhance schema representations and alignment between natural language and table content. These advances illustrate that pretrained models can surpass execution-guided seq2seq baselines. However, such results hold only for single-table WikiSQL and should be interpreted with caution when generalizing to multi-table or cross-domain settings
@@
Recognizing these limitations, new benchmarks were introduced. Spider [Spider2018] established a cross-domain standard with multi-table queries, JOINs, and nested structures, requiring compositional reasoning. UNITE [Lan2023UNITE] aggregated multiple datasets for broader SQL pattern coverage, while BIRD [Wretblad2024BIRD] emphasized robustness under noisy annotations. These resources expand scope, but none directly address the persistent quality issues in WikiSQL itself.
@@
A critical body of work addresses annotation quality, evaluation methodology, and robustness. Finegan-Dollak et al. [FineganDollak2018] analyzed evaluation practices and demonstrated that certain dataset splits and anonymization choices can inflate reported performance; they proposed more conservative splitting strategies and evaluation protocols to better assess compositional generalization. Recent research, such as studies on BIRD-Bench, has highlighted that crowd-sourced datasets often contain issues such as noisy annotations, mismatched data types, and questions that cannot be answered from the table. These problems can bias standard evaluation metrics, including both execution accuracy and logical-form accuracy, giving a misleading impression of model performance [Wretblad2024BIRD]. To address this, researchers recommend cleaning datasets and performing explicit verification of model outputs. Additionally, broader benchmark analyses emphasize the need for standardized evaluation procedures and the inclusion of realistic, user-generated queries, which help ensure that models are tested on scenarios closer to real-world usage and that reported metrics more accurately reflect practical performance [EDBT2025Analysis, Lan2023UNITE].
@@
Taken together, the literature highlights three concurrent directions: (1) architectural and schema-aware improvements to increase generation correctness (e.g., type-aware models, table-aware encoding models, execution-guided decoding) [Hwang2019SQLova, Yu2018TypeSQL, Wang2018EG]; (2) the design of more representative and cleaner benchmarks and evaluation practices that mitigate annotation noise (e.g., Spider, UNITE, and the recommendations of Finegan-Dollak) [Spider2018, FineganDollak2018, Lan2023UNITE]; and (3) the systematic evaluation of large pretrained and/or instruction-tuned LMs for Text-to-SQL, along with the development of practical engineering techniques (prompting, constrained decoding, schema linking, execution validation) required to obtain executable and reliable queries in real-world settings [Liu2024SurveyLLM, Zhang2024BenchmarkLLM].
@@ III-A Introduction to NL2SQL and the WikiSQL Dataset
WikiSQL contains 80,654 (question, SQL) pairs derived from 24,241 Wikipedia tables [Zhong2017Seq2SQL]. It has become a standard benchmark due to its scale, simple query structures, and public availability.
@@ IV-1 Temperature
Recent studies have shown that using greedy decoding to evaluate reasoning models can lead to higher repetition rates and significant variability across different checkpoints [deepseek_r1]. Therefore, for models with strong reasoning capabilities, we did not use greedy decoding. Instead, we employed the default sampling parameters recommended for each model.
@@ IV-4 SQL Extraction from Outputs
To address this, we adopted a regex-based extraction strategy similar to what is used in popular text-to-SQL benchmarks such as Spider [Spider2018]. Specifically, we designed a regex pattern to extract all possible SQL statements from the model output:
@@ IV-A Main Scenario (Benchmark)
We evaluated a broad set of models across different scenarios, including Llama 3.2 1B Instruct, Qwen 2.5 1.5B Instruct, Phi 3.5 Mini Instruct, Gemma 3 4B IT, Gemma 3 27B IT, Mistral 7B Instruct v0.2, gpt-oss-20B, DeepSeek V3 0324, DeepSeek R1 0528, and OpenAI o4-mini [llama3_paper, qwen2.5, phi3.5, gemma_2025, mistral_7b, gpt_oss, deepseek_v3, deepseek_r1, openai_o4_mini].
@@ V Conclusions
We also argue that, despite its simplicity, LLMSQL remains a relevant and informative benchmark. Real-world SQL workloads are often dominated by relatively simple patterns: in an analysis of 8.1 million production queries at Uber [johnson2018practical], over 62% used JOIN, and fewer than 1% involved operators such as UNION, INTERSECT, or EXCEPT. This reinforces the idea that mastering simple clauses is critical for practical deployment and benchmarking progress.